OpenStudy (anonymous):

Alg...Please help? What is the product of 2*sqrt(35x)times3*sqrt(14x) Simplify if possible...help..anyone?

OpenStudy (lgbasallote):

\[2\sqrt{35 x} \times 3\sqrt{14x} \implies (2\times 3)(\sqrt{35x} \times \sqrt{14x}) \implies (2\times 3)(\sqrt{35x \times 14x})\] does that help?

OpenStudy (anonymous):

Yes, yes it does...Thank you! ^_^